1.The purchasing power of the $20 bill increases over time due to inflation. Group of answer choices True False 2.The consumer price index increased from 120 to 132. If you received a raise equal to 10% during this time period then your real income has decreased. Group of answer choices True False 3.Productivity is measured by calculating the growth rate of real GDP. Group of answer choices True False 4.The labor force is made up of all people who are at least 16 years old and are capable of working. Group of answer choices True False 5.If more spouses decide to quit their jobs to stay home with their young children the supply of labor will ______ causing the average wage to _______. Group of answer choices increase : increase decrease : decrease increase : decrease decrease : increase 6.If your employer gives you a raise that is less than the inflation rate, then your real salary will have declined. Group of answer choices True False 7.An increase in the minimum wage is likely to _____ the quantity of labor demanded and at the same time is likely to _____ the quantity of labor supplied. Group of answer choices decrease : decrease increase : decrease increase : increase decrease : increase 8.To measure the change in the standard of living of the average person in a country, we most often calculate the percentage change in real GDP per capita. Group of answer choices True False 9.All else the same, as the baby boom generation (born 1946 - 1964) begins to retire, the rate of unemployment will decline. Group of answer choices True False 10.Countries with unemployment insurance programs are likely to have higher levels of unemployment. Group of answer choices True False 11.If the average wage paid to the worker was $20 in the year 1990 and $30 in the year 2000, then the average worker in the year 2000 must have been better off in terms of purchasing power. Group of answer choices True False 12.At an annual growth rate of 2.8% it will take years for a country’s GDP to double. If GDP starts at a value of $10 billion, then in 75 years we would expect the value of GDP to be times larger. Enter whole numbers. 13.A company produces 100 cars in the year 2017 but only manages to sell 90 of the cars. In the year 2018 they sell the 10 cars that were not sold in the previous year to consumers. In which spending categories of GDP do we record the 10 cars? Group of answer choices The unsold cars count as consumption spending in 2018. The unsold cars count in the inventory investment category for both years and count as consumption spending in 2018. The unsold cars count as inventory investment in 2017 and consumption spending in 2018. The unsold cars count as inventory investment for 2017. 2018 GDP is not affected. 14.The inflation rate between the years 2000 and 2001 was 4.18%. Based on this information, a basket of goods that cost $157 in the year 2000 would now cost how much in the year 2001? Enter a number rounded to two decimal places.
Answer #1

1. The correct option for this question will be false. Inflation always decreases the purchasing power of money. As increase in price level means that now with the same $20 bill you'll be able to buy less.

2. The correct option for this question will be false. Since the rise in inflation that is percentage change in CPI is also 10%.

Rate of inflation = 132 - 120/120 ×100

= 12/120 × 100 = 10%

The rate of inflation is 10% for given period which is equal to the raise in pay which is also 10%. So the the real income hasn't decreased.

3. The correct option for this question will be false. The growth rate of real GDP means we measure the growth rate of output at constant prices which means we allow for the factor of productions to increase but not prices. The growth rate of productivity means that factors of production doesn't change and yet output grows which is different from growth rate of real GDP.

4. The correct option for this question will be true. By definition the labor force is made of all the people who are above 16 and capable of working.

5. The correct option for this question will be, decrease increase. Beacuse when more spouse decides to stay at home the supply of labor will decrease and this will cause the average wage to increase.

6. The correct option for this question will be true. Because when raise is less than the inflation rate then the real purchasing power decreases.

7. The correct option for this question will be decrease increase. Beacuse the wage is the price that employer pays the employees thus when wage increases, employees demand less quantity of labor causing a labor demand to decrease. While at same time wage is price workers get for providing their labor to employer, thus increase in wage increase the quantity of labor supply.

8. The correct option for this question will be true. Real per capita gdp shows how the purchasing power of an average person has increased over time, which indeed decides the standard of living.

9. The correct option for this question will be true. As more people retire the labor force decreases. And unemployment rate is equal to = total unemployes people/ total labor force.

As you can see as the total labor force decreases the unemployment rate decreases.

10. The correct option for this question will be true.

11. The correct option for this question will be false. Because we only know the rise in average wage of person from year 1990 to 2000 but we know nothing about the inflation. If rise in inflation was more than rise in wage then purchasing power will decrease. So we can't say for sure that real wage has increased in terms of purchasing power.

12. With annual growth rate of 2.8% the countries GDP after 75 years will increase (1+0.028)^75 = 8.

The countries GDP will increase 8 times.

13. The correct option for this question will be the last option that is unsold cars are considered as inventory investment in 2017 and 2018 GDP is not affected.

14. The rate of inflation = (prices in 2001 - prices in 2000/ prices in 2000) × 100

= 4.8 = (prices in 2001 - 157/157) × 100

= 4.8/100 = (prices in 2001 - 157)/157

= 0.048 × 157 + 157 = prices in 2001

= 7.536 + 157 = prices in 2001

= $164.53 = prices in 2001.